# Tele-LLMs: A Series of Specialized Large Language Models for Telecommunications

### v2

@@ 3.3. Wikipedia
Through this two-step process, we curate a dataset of 19.5k technically relevant telecommunications articles from Wikipedia. Similarly, we evaluated our LLM-based filtering by sampling 500 flagged articles and manually annotating them. We then assessed the precision and recall of our filtering process. The results in Table 2 showcase the high recall of the process, showcasing its ability to identify relevant technical content in telecommunications. The moderate precision, on the other hand, stems from the challenge of determining the appropriate level of technicality required for an article, resulting in false positives within the dataset.
3.4. Websites
The last source of telecommunications material we consider is the Common Crawl dataset, which comprises web archives from across the internet. To avoid issues with duplicates, non-English content, and potential profanity found in raw dumps, we utilized the refined web dataset (Penedo et al., 2023). This curated version of Common Crawl contains approximately 1 billion rows across 2.8 terabytes of data. To further eliminate duplicates, we filtered out Wikipedia articles from the dataset. Next, to extract telecommunications-related content from this refined dataset, we employed the same two-step process used for Wikipedia articles. Additionally, we incorporated content from well-known telecommunications blogs, such as ShareTechNote, to enhance the dataset’s relevance. The resulting collection consists of content from 740k website links, providing a comprehensive representation of telecommunications information available on the web.
@@ 3.5. Dataset Format
4. Evaluation Dataset
After preparing Tele-Data, the next step involves creating an evaluation dataset to test the resulting domain-adapted models’ telecommunications knowledge. Currently, only one such dataset exists: TeleQnA (Maatouk et al., 2023). TeleQnA is a multiple-choice question (MCQ) dataset drawn from standards and research papers, generated with human involvement. Although an MCQ dataset simplifies evaluation in terms of accuracy, it remains limiting because LLMs are not robust MCQ selectors due to their inherent ‘selection bias,’ a bias prevalent in virtually all LLMs (Zheng et al., 2024; Griot et al., 2024).
@@ 4. Evaluation Dataset
To overcome this challenge, we adopt an LLM-based approach to create our evaluation dataset, which we refer to as Tele-Eval. Specifically, we gather material from papers, standards, and wiki articles within the collected Tele-Data. We focus on these categories due to their relative cleanliness compared to the Common Crawl data. We then segment the material into 20,000-character chunks. Using Mixtral 8x7B-Instruct, we prompt the LLM to generate five questions for each segment in a three-shot manner, providing examples similar to those in Trivia QA (Joshi et al., 2017), but tailored to the telecommunications domain. The generated questions then undergo extensive regular expression-based filtering. Particularly, we implement filters to remove questions containing expressions such as ‘in this context,’ ‘in this paper,’ ‘as highlighted in,’ and references to equations, figures, tables, and other local elements, along with 100 other similar filters. Our aim is to achieve the lowest possible retention rate to ensure the highest quality in the automatically generated dataset.
@@
Remark 2.
One important aspect of our dataset is that each generated QnA pair includes the ID of the specific telecommunications material upon which it was based. This allows users to select questions from the material they are interested in and facilitating retrieval-augmented generation frameworks (Lewis et al., 2021). For example, if someone is interested in a specific topic (e.g., source and channel coding), they can identify the relevant content in the telecommunications dataset, obtain their IDs, and then use the portion of Tele-Eval linked to those IDs.
@@
5. Evaluation Metrics
One of the key challenges after continual pretraining is effectively evaluating the resulting domain-adapted model. As seen in the previous section, Tele-Eval consists of open-ended question-answer pairs. Evaluating open-ended responses is more complex compared to MCQ datasets. For example, traditional evaluation metrics such as ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002), commonly used for summarization and translation tasks, measure lexical overlap between the model’s output and the reference answers. However, they fail to capture the semantic similarity and correctness of responses when the model uses alternative wording or lexicon to convey the same meaning as the reference answers. This challenge is especially pronounced when evaluating equations generated by the models, as is often required in the telecommunications domain, since these metrics struggle to capture such nuances. To address this, we adopted three evaluation metrics: Answer perplexity, SemScore (Aynetdinov and Akbik, 2024), and LLM-Eval (Zheng et al., 2023). In Section 7, we will showcase the pros and cons of each evaluation metric and conclude that LLM-Eval is the most robust comparative tool for the telecommunications domain. Below, we provide details on each of these metrics.
5.1. Answer Perplexity
@@
(5) $\begin{split}\text{Ans-PPL}(\bm{x})=\exp\bigg{(}&-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=k_{i}}^{T_{i}}\frac{1}{T_{i}-k_{i}+1}\\ &\log P(x^{i}_{j}\mid\bm{x}^{i}_{1:j-1};\bm{W})\bigg{)},\end{split}$
@@
5.2. SemScore
Another metric we use to evaluate the correctness of a model’s output relative to the ground truth answer is semantic similarity. This is achieved by leveraging sentence transformer models (Reimers and Gurevych, 2020), such as the all-mpnet-base model [^7] https://sbert.net/docs/sentence_transformer/pretrained_models.html from the Sentence Transformers family. These encoder-only BERT models are trained using contrastive loss to encode pairs of sequences into a high-dimensional space, such that the cosine similarity between embeddings of similar pairs is high (closer to 1), while it is lower for dissimilar ones (closer to -1). With this in mind, we define the SemScore (Aynetdinov and Akbik, 2024) as
(6) $\textnormal{SemScore}(\bm{x},\bm{y})=\cos(\textnormal{embed}(\bm{x}),\textnormal{embed}(\bm{y})),$
@@ 5.2. SemScore
5.3. LLM-Eval
Figure 4. Training metrics for Gemma-2B (top) and Llama-3-8B (bottom) models using LoRa.
The last metric in our framework involves using an LLM as a judge to assess the correctness of a model’s output compared to the ground truth answer (Zheng et al., 2023). Given the extensive knowledge LLMs have acquired through their training, we can leverage them as assistants to evaluate the quality and accuracy of model outputs. Previous studies have demonstrated that LLM-based judges can effectively match both controlled and crowdsourced human preferences in various evaluation tasks (Zheng et al., 2023). With this approach, when evaluating any model outputs for questions in Tele-Eval, we prompt Mixtral 8x-7B-Instruct to compare the model’s output to the ground truth and provide a Yes or No answer regarding its correctness. We define the LLM-Eval metric as the average score of these boolean responses, such that
@@ 5.3. LLM-Eval
6. Initial Experiments
With both the training and evaluation datasets prepared, the next step is to train our models. However, before proceeding, several key questions regarding training techniques and parameters need to be addressed to lay the foundation for the training procedure. While general guidelines on training parameters, such as learning rate and batch size, can be drawn from previous research (Labrak et al., 2024; Colombo et al., 2024; Xie et al., 2023), the characteristics of telecommunications knowledge—such as the prevalence of equations—introduce specific challenges. Key questions that need to be answered include whether full parameter fine-tuning is necessary or if parameter-efficient tuning methods are sufficient, as well as determining the number of training epochs needed to effectively adapt the LLM to the telecommunications domain. We address these questions in this section, while noting that our results can serve as a guide for adapting other models to the telecommunications domain beyond those discussed in our paper.
6.1. Training Settings
For these experiments, we consider two models: Gemma-2B and LLaMA-3-8B. Throughout these studies, we use 20% of the Tele-Data dataset as the training data, with 10% of this subset serving as the validation set. To solve eq. (2), we set the batch size to 4M tokens (with a sequence length of 8192 tokens) and use the AdamW optimizer with a weight decay of 0.1, keeping the default Hugging Face $\beta$ parameters unchanged. The maximum gradient norm is set to $1$. The learning rate is decreased according to a cosine learning rate schedule, down to 10% of the maximum learning rate, and a linear warmup of 10% of an epoch is applied (Gupta et al., 2023). The maximum learning rate is set to 1e-5 for LLaMA and 2e-5 for Gemma. These values are aligned with the pretraining stages of these LLMs (et al., 2024; Team, 2024). To improve efficiency, we employ mixed-precision training alongside sample packing and block attention techniques to avoid cross-contamination (Krell et al., 2022).
6.2. PEFT vs. FFT
PEFT techniques aim to reduce computational costs and memory requirements during domain adaptation by training only a minimal number of parameters, in contrast to FFT methods that update all the model’s parameters. Among these techniques, LoRa (Hu et al., 2022) stands out as a prominent PEFT method. It adds small, trainable low-rank matrices to the existing weights of the pretrained model. To investigate whether PEFT methods like LoRa are sufficient for adapting language models to the telecommunications domain, we conducted a comparative study between LoRa and FFT using both Gemma-2B and LLaMA-3-8B models. For our LoRa implementations, we fixed the LoRa rank to r=64, set the Lora alpha to 32, and used a LoRa dropout rate of 0.1.
@@ 6.1. Training Settings
6.3. One or Multiple Epochs?
Figure 5. Training metrics for Gemma-2B (top) and Llama-3-8B (bottom) models across three epochs.
@@ 6.3. One or Multiple Epochs?
7. Tele-LLMs
Figure 6. Training loss of the Tele-LLMs series.
@@ 7. Tele-LLMs
Table 3. LLMs’ performance across Tele-Eval and other general knowledge benchmarks.
@@
Having completed the initial experiments, we proceed with the full-scale training of the LLMs to create the telecom-adapted series of models, starting with the following three models: Tinyllama-1.1B-Tele, Gemma-2B-Tele, and LLaMA-3-8B-Tele. Our selection of these base models was guided by a combination of factors, including their performance on the Hugging Face Open LLM leaderboard and a consideration of their licenses. We also accounted for the diversity of the releasing companies to mitigate potential risks associated with changes in these licenses.
@@
Training Settings. We maintain the settings described in Section 6.1, with Tinyllama-1.1B following the same approach as Gemma-2B. One caveat is the use of a sequence length of 2048 for Tinyllama-1.1B, aligning with its initial pretraining specifications (et al., 2023). For the training dataset across the entire series, we utilize the entire Tele-Data dataset for two epochs, augmented with 5% of regular general-purpose data taken from SlimPajama [^8] https://huggingface.co/datasets/cerebras/SlimPajama-627B. This augmentation is designed to mitigate the catastrophic forgetting phenomenon by reintroducing general-purpose data to the model, helping it retain key skills acquired during the initial pretraining stage. The training of these models was conducted on a cluster of 8 NVIDIA A6000 GPUs, consuming a total of 2,500 GPU hours.
@@
Loss Trends. As shown in Fig. 6, the three models exhibit results consistent with those reported in our previous section’s training experiments. Specifically, Tinyllama-1.1B and Gemma-2B demonstrate the steady training loss behavior typical of smaller models, while LLaMA-3-8B displays the sharp loss reduction characteristic of larger models at each epoch. On another note, due to its shorter context length, Tinyllama-1.1B starts with a lower training loss than Gemma-2B, highlighting the challenge in capturing long-range dependencies within 8k token sequences.
@@
Evaluation Settings. We evaluate the LLMs on Tele-Eval using the three metrics discussed in Section 5, generating model answers through greedy decoding of 100 new tokens with the prompt provided in Appendix C. To assess the impact of our training on other LLM capabilities, we evaluate our models on three additional benchmarks: MMLU (Hendrycks et al., 2021) for general world knowledge, HellaSWAG (Zellers et al., 2019) for commonsense understanding, and GSM-8K (Cobbe et al., 2021) for mathematical reasoning abilities. All evaluations, including Tele-Eval and the additional benchmarks, are conducted in a zero-shot setting.
@@
Quantitative Evaluation. As shown in Table 3, the adapted telecom LLMs outperform their counterparts across all metrics on Tele-Eval. For general knowledge tasks, these models exhibit stable performance compared to the original models, with slight variations—either improvements or degradations—across most tasks. Notably, there are improvements on GSM-8K, with LLama-3-8B-Tele achieving a 10% higher accuracy. This enhancement is attributed to the equations found in Tele-Data, which enable the adapted model to become stronger in mathematical reasoning.
@@
Qualitative Evaluations. Going beyond the quantitative comparison provided in Table 3, we offer a set of examples in Appendix D that illustrate how the behavior of these models change, using Gemma-2B and Gemma-2B-Tele as a case study. These examples highlight the model’s enhanced ability to provide detailed explanations, confidently answer telecommunications-related questions, and even delve into telecommunications concepts when they are mentioned in the prompt. These qualitative observations further emphasize the successful adaptation of these models to the telecommunications domain.
8. Further Exploration
@@ 8.1. Division of Expertise
Table 4. Performance comparison between Gemma-2B, Gemma-2B-Tele, and Gemma-2B-Standards.
When adapting an LLM to the telecommunications domain, a key question arises: should we adapt a single LLM to cover the entire domain, or should we adapt multiple LLMs, each specialized in a specific aspect of telecommunications? For example, one LLM could focus on scholarly material while another concentrates on standards. To explore this, we compare two versions of the Gemma-2B model: one trained exclusively on the standards portion of Tele-Data, referred to as Gemma-2B-Standards, and Gemma-2B-Tele. The training parameters align with those previously reported for Gemma-2B-Tele. Evaluation is based on both the entirety of Tele-Eval and questions related to standards using the LLM-eval metric, along with general knowledge evaluation datasets to assess the retention of the model’s capabilities. The results are reported in Table 4.
@@
8.2. Pretraining Data Impact

### v3

@@ III-C Wikipedia
Through this two-step process, we curate a dataset of 19.5k technically relevant telecommunications articles from Wikipedia. Similarly, we evaluated our LLM-based filtering by sampling 500 flagged articles and manually annotating them. We then assessed the precision and recall of our filtering process. The results in Table II showcase the high recall of the process, showcasing its ability to identify relevant technical content in telecommunications. The moderate precision, on the other hand, stems from the challenge of determining the appropriate level of technicality required for an article, resulting in false positives within the dataset.
III-D Websites
The last source of telecommunications material we consider is the Common Crawl dataset, which comprises web archives from across the internet. To avoid issues with duplicates, non-English content, and potential profanity found in raw dumps, we utilized the refined web dataset [21]. This curated version of Common Crawl contains approximately 1 billion rows across 2.8 terabytes of data. To further eliminate duplicates, we filtered out Wikipedia articles from the dataset. Next, to extract telecommunications-related content from this refined dataset, we employed the same two-step process used for Wikipedia articles. Additionally, we incorporated content from well-known telecommunications blogs, such as ShareTechNote, to enhance the dataset’s relevance. The resulting collection consists of content from 740k website links, providing a comprehensive representation of telecommunications information available on the web.
@@ III-E Dataset Format
IV Evaluation Dataset
After preparing Tele-Data, the next step involves creating an evaluation dataset to test the resulting domain-adapted models’ telecommunications knowledge. Currently, only one such dataset exists: TeleQnA [22]. TeleQnA is a multiple-choice question (MCQ) dataset drawn from standards and research papers, generated with human involvement. Although an MCQ dataset simplifies evaluation in terms of accuracy, it remains limiting because LLMs are not robust MCQ selectors due to their inherent ‘selection bias,’ a bias prevalent in virtually all LLMs [23, 24].
@@ IV Evaluation Dataset
To overcome this challenge, we adopt an LLM-based approach to create our evaluation dataset, which we refer to as Tele-Eval. Specifically, we gather material from papers, standards, and wiki articles within the collected Tele-Data. We focus on these categories due to their relative cleanliness compared to the Common Crawl data. We then segment the material into 20,000-character chunks. Using Mixtral 8x7B-Instruct, we prompt the LLM to generate five questions for each segment in a three-shot manner, providing examples similar to those in Trivia QA [25], but tailored to the telecommunications domain. The generated questions then undergo extensive regular expression-based filtering. Particularly, we implement filters to remove questions containing expressions such as ‘in this context,’ ‘in this paper,’ ‘as highlighted in,’ and references to equations, figures, tables, and other local elements, along with 100 other similar filters. Our aim is to achieve the lowest possible retention rate to ensure the highest quality in the automatically generated dataset.
@@
Remark 2
One important aspect of our dataset is that each generated QnA pair includes the ID of the specific telecommunications material upon which it was based. This allows users to select questions from the material they are interested in and facilitating retrieval-augmented generation frameworks [18]. For example, if someone is interested in a specific topic (e.g., source and channel coding), they can identify the relevant content in the telecommunications dataset, obtain their IDs, and then use the portion of Tele-Eval linked to those IDs.
@@
V Evaluation Metrics
One of the key challenges after continual pretraining is effectively evaluating the resulting domain-adapted model. As seen in the previous section, Tele-Eval consists of open-ended question-answer pairs. Evaluating open-ended responses is more complex compared to MCQ datasets. For example, traditional evaluation metrics such as ROUGE [26] and BLEU [27], commonly used for summarization and translation tasks, measure lexical overlap between the model’s output and the reference answers. However, they fail to capture the semantic similarity and correctness of responses when the model uses alternative wording or lexicon to convey the same meaning as the reference answers. This challenge is especially pronounced when evaluating equations generated by the models, as is often required in the telecommunications domain, since these metrics struggle to capture such nuances. To address this, we adopted three evaluation metrics: Answer perplexity, SemScore [28], and LLM-Eval [29]. In Section VII, we will showcase the pros and cons of each evaluation metric and conclude that LLM-Eval is the most robust comparative tool for the telecommunications domain. Below, we provide details on each of these metrics.
V-A Answer Perplexity
@@
$\begin{split}\text{Ans-PPL}(\bm{x})=\exp\bigg{(}&-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=k_{i}}^{T_{i}}\frac{1}{T_{i}-k_{i}+1}\\ &\log P(x^{i}_{j}\mid\bm{x}^{i}_{1:j-1};\bm{W})\bigg{)},\end{split}$ (5)
@@
V-B SemScore
Another metric we use to evaluate the correctness of a model’s output relative to the ground truth answer is semantic similarity. This is achieved by leveraging sentence transformer models [30], such as the all-mpnet-base model [^7] https://sbert.net/docs/sentence_transformer/pretrained_models.html from the Sentence Transformers family. These encoder-only BERT models are trained using contrastive loss to encode pairs of sequences into a high-dimensional space, such that the cosine similarity between embeddings of similar pairs is high (closer to 1), while it is lower for dissimilar ones (closer to -1). With this in mind, we define the SemScore [28] as
$\textnormal{SemScore}(\bm{x},\bm{y})=\cos(\textnormal{embed}(\bm{x}),\textnormal{embed}(\bm{y})),$ (6)
@@ V-B SemScore
V-C LLM-Eval
Figure 4: Training metrics for Gemma-2B (top) and Llama-3-8B (bottom) models using LoRa.
The last metric in our framework involves using an LLM as a judge to assess the correctness of a model’s output compared to the ground truth answer [29]. Given the extensive knowledge LLMs have acquired through their training, we can leverage them as assistants to evaluate the quality and accuracy of model outputs. Previous studies have demonstrated that LLM-based judges can effectively match both controlled and crowdsourced human preferences in various evaluation tasks [29]. With this approach, when evaluating any model outputs for questions in Tele-Eval, we prompt Mixtral 8x-7B-Instruct to compare the model’s output to the ground truth and provide a Yes or No answer regarding its correctness. We define the LLM-Eval metric as the average score of these boolean responses, such that
@@ V-C LLM-Eval
VI Initial Experiments
With both the training and evaluation datasets prepared, the next step is to train our models. However, before proceeding, several key questions regarding training techniques and parameters need to be addressed to lay the foundation for the training procedure. While general guidelines on training parameters, such as learning rate and batch size, can be drawn from previous research [13, 14, 15], the characteristics of telecommunications knowledge—such as the prevalence of equations—introduce specific challenges. Key questions that need to be answered include whether full parameter fine-tuning is necessary or if parameter-efficient tuning methods are sufficient, as well as determining the number of training epochs needed to effectively adapt the LLM to the telecommunications domain. We address these questions in this section, while noting that our results can serve as a guide for adapting other models to the telecommunications domain beyond those discussed in our paper.
VI-A Training Settings
For these experiments, we considered two models: Gemma-2B and LLaMA-3-8B. This choice was motivated by our intention to explore the behavior of model training across different parameter sizes. Throughout these studies, we use 20% of the Tele-Data dataset as the training data, with 10% of this subset serving as the validation set. To solve eq. (2), we set the batch size to 4M tokens (with a sequence length of 8192 tokens) and use the AdamW optimizer with a weight decay of 0.1, keeping the default Hugging Face $\beta$ parameters unchanged. The maximum gradient norm is set to $1$. The learning rate is decreased according to a cosine learning rate schedule, down to 10% of the maximum learning rate, and a linear warmup of 10% of an epoch is applied [31]. The maximum learning rate is set to 1e-5 for LLaMA and 2e-5 for Gemma. These values are aligned with the pretraining stages of these LLMs [3, 32]. To improve efficiency, we employ mixed-precision training alongside sample packing and block attention techniques to avoid cross-contamination [33].
VI-B PEFT vs. FFT
Figure 5: Training metrics for Gemma-2B (top) and Llama-3-8B (bottom) models across three epochs.
PEFT techniques aim to reduce computational costs and memory requirements during domain adaptation by training only a minimal number of parameters, in contrast to FFT methods that update all the model’s parameters. Among these techniques, LoRa [34] stands out as a prominent PEFT method. It adds small, trainable low-rank matrices to the existing weights of the pretrained model. To investigate whether PEFT methods like LoRa are sufficient for adapting language models to the telecommunications domain, we conducted a comparative study between LoRa and FFT using both Gemma-2B and LLaMA-3-8B models. For our LoRa implementations, we fixed the LoRa rank to r=64, set the Lora alpha to 32, and used a LoRa dropout rate of 0.1.
@@ VI-A Training Settings
VI-C One or Multiple Epochs?
@@
VII Tele-LLMs
Figure 6: Training loss of the Tele-LLMs series.
@@ VII Tele-LLMs
Gemma-2-2B 12.96 0.5853 15.10 0.4789 0.5164 0.0607
Gemma-2-2B-Tele 10.86 0.6237 17.34 0.4773 0.5212 0.0697
LLama-3.2-1B 11.72 0.6205 11.85 0.3611 0.4769 0.0288
LLama-3.2-1B-Tele 10.50 0.6318 12.98 0.3631 0.4751 0.0303
LLama-3.2-3B 9.52 0.6413 24.68 0.5409 0.5522 0.1152
LLama-3.2-3B-Tele 9.16 0.6557 26.71 0.5366 0.5489 0.1008
@@
TABLE III: LLMs’ performance across Tele-Eval and other general knowledge benchmarks.
@@
Having completed the initial experiments, we proceed with the full-scale training of the LLMs to create the telecom-adapted series of models, starting with the following models: TinyLlama-1.1B, Gemma-2B, Gemma-2-2B, LLaMA-3.2-1B, LLaMA-3.2-3B, and LLaMA-3-8B. Our selection of these base models was guided by a combination of factors, including their performance on the Hugging Face Open LLM leaderboard and a consideration of their licenses. We also accounted for the diversity of the releasing companies to mitigate potential risks associated with changes in these licenses.
@@
Training Settings. We maintain the training settings described in Section VI-A, where models smaller than 8B parameters follow the same approach as Gemma-2B while aligning the sequence length to each model’s pretraining settings. For the training dataset across the entire series, we utilize the entire Tele-Data dataset for two epochs, augmented with 5% of regular general-purpose data taken from SlimPajama [^8] https://huggingface.co/datasets/cerebras/SlimPajama-627B. This augmentation is designed to mitigate the catastrophic forgetting phenomenon by reintroducing general-purpose data to the model, helping it retain key skills acquired during the initial pretraining stage. The training of these models was conducted on a cluster of 8 NVIDIA A6000 GPUs, consuming a total of 5,000 GPU hours.
@@
Loss Trends. In Fig. 6, we report the training loss across three models. Note that all sub-8B parameter models exhibited similar trends to those of Gemma-2B and Tinyllama-1.1B shown in Fig. 6. These models, in particular, demonstrate steady training loss behavior, while LLaMA-3-8B displays the sharp loss reduction characteristic of larger models at each epoch depicted in the previous section.
@@
Evaluation Settings. We evaluate the LLMs on Tele-Eval using the three metrics discussed in Section V, generating model answers through greedy decoding of 100 new tokens with the prompt provided in Appendix C. To assess the impact of our training on other LLM capabilities, we evaluate our models on three additional benchmarks: MMLU [35] for general world knowledge, HellaSWAG [36] for commonsense understanding, and GSM-8K [37] for mathematical reasoning abilities. All evaluations, including Tele-Eval and the additional benchmarks, are conducted in a zero-shot setting.
@@
Quantitative Evaluation. As shown in Table III, the adapted telecom LLMs outperform their counterparts across all metrics on Tele-Eval. For general knowledge tasks, these models exhibit stable performance compared to the original models, with slight variations—either improvements or degradations—across most tasks. Notably, there are improvements on GSM-8K, with LLama-3-8B-Tele achieving a 10% higher accuracy. This enhancement is attributed to the equations found in Tele-Data, which enable the adapted model to become stronger in mathematical reasoning.
@@
Qualitative Evaluations. Going beyond the quantitative comparison provided in Table III, we offer a set of examples in Appendix D that illustrate how the behavior of these models change, using Gemma-2B and Gemma-2B-Tele as a case study. These examples highlight the model’s enhanced ability to provide detailed explanations, confidently answer telecommunications-related questions, and even delve into telecommunications concepts when they are mentioned in the prompt. These qualitative observations further emphasize the successful adaptation of these models to the telecommunications domain.
VIII Further Exploration
@@ VIII-A Division of Expertise
TABLE IV: Performance comparison between Gemma-2B, Gemma-2B-Tele, and Gemma-2B-Standards.
When adapting an LLM to the telecommunications domain, a key question arises: should we adapt a single LLM to cover the entire domain, or should we adapt multiple LLMs, each specialized in a specific aspect of telecommunications? For example, one LLM could focus on scholarly material while another concentrates on standards. To explore this, we compare two versions of the Gemma-2B model: one trained exclusively on the standards portion of Tele-Data, referred to as Gemma-2B-Standards, and Gemma-2B-Tele. The training parameters align with those previously reported for Gemma-2B-Tele. Evaluation is based on both the entirety of Tele-Eval and questions related to standards using the LLM-eval metric, along with general knowledge evaluation datasets to assess the retention of the model’s capabilities. The results are reported in Table IV.
@@
VIII-B Pretraining Data Impact